### v1

@@ 6 Conclusions
Firm-to-firm relationships are a central feature of international trade. These relationships often involve market power on both sides, with prices determined through bilateral negotiations rather than market-clearing conditions. We develop a framework that departs from standard models of price-taking buyers by allowing both importers and exporters to influence price formation through bargaining. This approach yields analytical and empirical tools to study how market power and network frictions shape prices, markups, and the transmission of shocks in firm-to-firm trade.
Using transaction-level data, we show that U.S. importers wield substantial bargaining power and face upward-sloping residual supply, consistent with oligopsonistic behavior. Within ongoing firm-to-firm relationships, tariff pass-through during the 2018 trade war was incomplete, with exporters absorbing much of the tariff through cost-side adjustments. This muted price response is primarily driven by cost-side adjustments: while markups do adjust, their contribution to pass-through is limited because strong buyer power endogenously flattens the markup elasticity. In this sense, the limited role of markups in shaping tariff responses is itself a consequence of strong importer power.
Although not the primary focus of our analysis, the results highlight the importance of bargaining for understanding the allocative implications of market power in firm-to-firm trade. In our model, when supplier power dominates, markup dispersion reflects inefficiencies and underproduction. By contrast, when buyer power is strong, the same dispersion can lead to efficient reallocation and higher aggregate output. These findings suggest that markup heterogeneity does not uniformly signal misallocation, and that its welfare consequences depend critically on the distribution of bargaining power. Extending the framework to study these welfare implications more formally remains a promising avenue for future work.
More broadly, our study offers a foundation for analyzing how market power shapes the incidence and transmission of shocks through the price channel. While we focus on detailed price-setting within firm-to-firm relationships, Acemoglu and Tahbaz-Salehi (2025) emphasize the general equilibrium implications of market power in production networks in shaping aggregate fluctuations. A promising direction for future research is to integrate these perspectives by combining micro-level bargaining dynamics with macro-level spillovers to study the broader implications of buyer and supplier power, including in domestic supply chains.
@@ Appendix A Mathematical Derivations
This appendix derives the key equations from the main text: the bilateral markup $\mu_{ij}$ (equation (2.6)) and the bilateral pass-through elasticity $\Phi_{ij}$ (equation (2.11)).
@@ A.1 Proof of Proposition 1
$q_{j}^{f}$ $=\gamma c_{j}q_{j}\left(p_{j}^{f}\right)^{-1},$ (A.1)
$q_{j}^{d}$ $=(\varrho-\gamma)c_{j}q_{j}\left(p_{j}^{d}\right)^{-1},$ (A.2)
where $c_{j}$ is the unit cost of output $q_{j}$, given by:
$c_{j}=\left[\varphi_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{\varrho-\gamma}\right]^{\frac{1}{\varrho}}q_{j}^{\frac{1-\varrho}{\varrho}}\equiv k_{j}q_{j}^{\frac{1-\varrho}{\varrho}}.$ (A.3)
@@
A relevant object for our derivations will be $\frac{d\ln c_{j}}{d\ln p_{j}^{f}},$ namely, the elasticity of the marginal cost $c_{j}$ with respect to $p_{j}^{f}.$ To find this elasticity, we first use the demand function downstream to write $q_{j}$ as
$q_{j}=\left(\frac{\nu}{\nu-1}\right)^{-\nu}c_{j}^{-\nu}D_{j},$ (A.4)
@@
$c_{j}$ $=\left[z_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{(\varrho-\gamma)}\right]^{\frac{1}{\varrho+\nu-\nu\varrho}}\left(\frac{\nu}{\nu-1}\right)^{-\nu\frac{1-\varrho}{\varrho+\nu-\nu\varrho}}\left(D_{j}\right)^{\frac{1-\varrho}{\varrho+\nu-\nu\varrho}},$
@@
$\frac{d\ln c_{j}}{d\ln p_{j}^{f}}=\frac{\gamma}{\varrho+\nu-\nu\varrho}.$
@@
The problem of the $i-j$ pair is to choose a bilateral price $p_{ij}$ that solves the following problem:
@@
00 0 $=\frac{d\ln\pi_{i}}{d\ln p_{ij}}+\frac{\phi}{1-\phi}\cdot\frac{GFT_{ij}^{i}}{\pi_{i}}\frac{\pi_{j}}{GFT_{ij}^{j}}\cdot\frac{d\ln\pi_{j}}{d\ln p_{ij}},$ (A.7)
@@
$\pi_{i}=p_{ij}q_{ij}+\sum_{k\neq j}p_{ik}q_{ik}-\theta c_{i}q_{i},$
where $c_{i}=k_{i}q_{i}^{\frac{1-\theta}{\theta}}.$ The elasticity of the profit $\pi_{i}$ with respect to $p_{ij}$ can be found as:
$\frac{d\ln\pi_{i}}{d\ln p_{ij}}$ $=\frac{q_{ij}}{\pi_{i}}(\varepsilon_{ij}-1)\left(\mu_{ij}^{\text{oligopoly}}c_{i}-p_{ij}\right),$ (A.8)
@@
$\tilde{\pi}_{i\left(-j\right)}=\sum_{k\neq j}p_{ik}q_{ik}-\theta\widetilde{c_{i}q_{i}}.$
@@
$\theta\widetilde{c_{i}q_{i}}$ $=\theta c_{i}q_{i}\left(\frac{\widetilde{q_{i}}}{q_{i}}\right)^{\frac{1}{\theta}}.$
$=\theta c_{i}q_{i}\left(1-x_{ij}\right)^{\frac{1}{\theta}}.$
@@
$=q_{ij}\left(p_{ij}-c_{i}\mu_{ij}^{\text{oligopsony}}\right),$
@@
$\pi_{j}=p_{j}q_{j}-\varrho c_{j}q_{j}=\frac{\gamma}{\eta-1}c_{j}q_{j},$
where $c_{j}$ is defined in equation (A.3), and $q_{j}$ in equation (A.4). The elasticity of the profit $\pi_{j}$ with respect to $p_{ij}$ can be found as:
$-\frac{d\ln\pi_{j}}{d\ln p_{ij}}=(\eta-1)s_{ij}.$ (A.10)
@@
$\tilde{\pi}_{j\left(-i\right)}=\frac{\gamma}{\eta-1}\widetilde{c_{j}q_{j}},$
where $\varrho\widetilde{c_{j}q_{j}}$ denotes the total “outside” cost of importer $j$. The assumptions on technology and demand downstream imply that we can write:
$\widetilde{c_{j}q_{j}}$ $=\left(1-s_{ij}\right)^{\frac{\eta-1}{\rho-1}}c_{j}q_{j}.$
@@
$\mu_{ij}:=\frac{p_{ij}}{c_{i}}=(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}+\omega_{ij}\mu_{ij}^{\text{oligopsony}}$ (A.12)
@@ A.2 Proof of Proposition 2
Proposition 2 characterizes how the bilateral markup $\mu_{ij}$ co-moves with the exporter’s and importer’s bilateral market shares, $s_{ij}$ and $x_{ij}$.
Part (i): $\mu_{ij}$ and Exporter’s supplier share $s_{ij}$
To study how $\mu_{ij}$ responds to the exporter’s supplier share $s_{ij}$, note that the oligopoly markup $\mu^{\text{oligopoly}}_{ij}$ is strictly increasing in $s_{ij}$ under standard assumptions ($\eta<\rho<\infty$). Differentiating the overall markup $\mu_{ij}$ in equation (A.12) with respect to $s_{ij}$ yields:
@@
$\frac{\partial\mu_{ij}}{\partial s_{ij}}=(1-\omega_{ij})\cdot\underbrace{\frac{\partial\mu^{\text{oligopoly}}_{ij}}{\partial s_{ij}}}_{>0}+\frac{\partial\omega_{ij}}{\partial s_{ij}}\cdot\underbrace{\left(\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}\right)}_{<0}.$
The first term is strictly positive when $\omega_{ij}<1$ (i.e., $\phi<1$), reflecting the direct effect of $s_{ij}$ on the oligopoly markup. The second term captures how changes in $s_{ij}$ affect the bargaining weight $\omega_{ij}$ and thus the relative influence of the oligopsony markdown. Although $\omega_{ij}$ is hump-shaped in $s_{ij}$, increasing at low values and decreasing at high values, its sensitivity to $s_{ij}$ is limited in most of the parameter space. As a result, $\frac{\partial\omega_{ij}}{\partial s_{ij}}$ is typically small, and the first (positive) term generally dominates.
In contrast, when $\phi=1$ (full importer bargaining power), the bilateral markup reduces to the oligopsony markdown: $\mu_{ij}=\mu^{\text{oligopsony}}_{ij}$, and $\omega_{ij}=1$. In this case,
$\frac{\partial\mu_{ij}}{\partial s_{ij}}=0.$
Thus, we have:
$\frac{\partial\mu_{ij}}{\partial s_{ij}}=\begin{cases}\gtrless 0&\text{if }\phi<1\text{ (typically positive)},\\ =0&\text{if }\phi=1.\end{cases}$
Thus, $\mu_{ij}$ can increase with $s_{ij}$ in our theory only when $\phi<1$, which proves the first part of Proposition 1.
Part (ii): $\mu_{ij}$ and Importer’s buyer share $x_{ij}$
The effect of $x_{ij}$ on $\mu_{ij}$ is limited to the oligopsony markdown term and is therefore easier to characterize. Differentiating the bilateral markup with respect to $x_{ij}$ yields:
$\frac{\partial\mu_{ij}}{\partial x_{ij}}=\omega_{ij}\cdot\underbrace{\frac{\partial\mu^{\text{oligopsony}}_{ij}}{\partial x_{ij}}}_{<0\ \text{if}\ \theta<1}.$
The markdown $\mu^{\text{oligopsony}}_{ij}$ decreases with $x_{ij}$ when marginal cost is increasing (i.e., $\theta<1$), as larger buyers elicit stronger cost reductions and thus negotiate lower prices. This implies:
$\frac{\partial\mu_{ij}}{\partial x_{ij}}<0\iff\omega_{ij}>0\text{ and }\theta<1.$
Therefore, observing a negative slope of $\mu_{ij}$ with respect to the importer’s share $x_{ij}$ provides evidence of both importer bargaining power ($\phi>0$ so that $\omega_{ij}>0$) and decreasing returns to scale ($\theta<1$). This completes the proof of Proposition 2. ∎
@@ A.3 Proof of Proposition 3
$\ln p_{ij}=\ln\mu_{ij}+\ln c_{i}+\ln T_{c},$
where $c_{i}$ and $\mu_{ij}$ are as in equations (2.1) and (2.6), respectively.
@@
where $\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}$ and $\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ are the partial markup and cost elasticities, respectively.
@@
$\ln c_{i}$ $=\ln k_{i}+\frac{1-\theta}{\theta}\ln q_{i}.$
@@
$\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ $=\frac{1-\theta}{\theta}\frac{d\ln q_{i}}{d\ln q_{ij}}\left(-\frac{d\ln q_{ij}}{d\ln p_{ij}}\right)$
@@
[…]
@@
[…]

### v2

@@ 6 Conclusions
We develop and estimate a structural model of two sided market power in firm-to-firm trade to shed light on tariff pass-through in global supply chains. The framework features oligopoly power on the exporter side and oligopsony power on the importer side, and yields closed-form expressions for bilateral markups and pass-through as functions of two sufficient statistics: the exporter’s supplier share and the importer’s buyer share, facilitating the mapping from theory to data, structural estimation, and the decomposition of pass-through mechanisms.
The two primitives governing oligopsony power, namely the inverse export supply elasticity and bargaining power, also govern the determinants of pass-through: the inverse export supply elasticity determines whether pass-through reflects cost or markup adjustments, while bargaining power shapes the nature of strategic markup adjustment. The framework therefore nests as special cases both the competitive benchmark under decreasing returns and the standard oligopoly pass-through model under constant returns.
Applied to U.S. imports in the 2018 tariff episode, our estimates reveal substantial oligopsony power and steep export supply curves. At these parameters, the cost channel dominates, explaining why pass-through within continuing buyer-supplier relationships is substantially incomplete and declines systematically with buyer share. This specific result likely reflects the relationship specificity and supply-side frictions that characterize U.S. firm-to-firm trade. In other settings, with weaker relationship specificity or flatter export supply, oligopsony power is lower and the framework accommodates a larger role for strategic markup adjustment.
Incomplete pass-through through the cost channel is thus a feature of short-run pricing in ongoing firm-to-firm relationships. Over longer horizons, as firms reallocate across partners and capacity constraints relax, effective returns to scale rise toward constant and the markup channel prevails, in line with the firm-level pass-through literature that attributes incomplete pass-through to strategic markup adjustment (e.g., amiti2014importers). Even there, our framework provides a richer characterization of how bargaining power shapes pass-through across bilateral relationships. Sample composition matters too: product-level studies pool repeated arm’s-length links with one-off and newly formed relationships, and we show that this extensive-margin variation accounts for the bulk of the near-complete pass-through documented at the product level (amiti2019impact; fajgelbaum2020return).
Our analysis is explicitly partial equilibrium. While we show that this approach yields accurate predictions of pass-through within firm-to-firm relationships, a full account of tariff incidence requires accounting for extensive-margin adjustments in the set of active relationships, as well as general-equilibrium responses in wages and export prices. Whether the cost-channel dominance we document at the intensive margin survives once these forces are taken into account is a key open question for future research.
@@ A.1 Proof of Proposition 1
$q_{j}^{f}$ $=\gamma MC_{j}q_{j}\left(p_{j}^{f}\right)^{-1},$ (A.1)
$q_{j}^{d}$ $=(\varrho-\gamma)MC_{j}q_{j}\left(p_{j}^{d}\right)^{-1},$ (A.2)
where $MC_{j}$ is the unit cost of output $q_{j}$, given by:
$MC_{j}=\left[\varphi_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{\varrho-\gamma}\right]^{\frac{1}{\varrho}}q_{j}^{\frac{1-\varrho}{\varrho}}\equiv k_{j}q_{j}^{\frac{1-\varrho}{\varrho}}.$ (A.3)
@@
A relevant object for our derivations will be $\frac{d\ln MC_{j}}{d\ln p_{j}^{f}},$ namely, the elasticity of the marginal cost $MC_{j}$ with respect to $p_{j}^{f}.$ To find this elasticity, we first use the demand function downstream to write $q_{j}$ as
$q_{j}=\left(\frac{\nu}{\nu-1}\right)^{-\nu}MC_{j}^{-\nu}D_{j},$ (A.4)
@@
$MC_{j}$ $=\left[z_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{(\varrho-\gamma)}\right]^{\frac{1}{\varrho+\nu-\nu\varrho}}\left(\frac{\nu}{\nu-1}\right)^{-\nu\frac{1-\varrho}{\varrho+\nu-\nu\varrho}}\left(D_{j}\right)^{\frac{1-\varrho}{\varrho+\nu-\nu\varrho}},$
@@
$\frac{d\ln MC_{j}}{d\ln p_{j}^{f}}=\frac{\gamma}{\varrho+\nu-\nu\varrho}.$
@@
The problem of the iji-j pair is to choose a bilateral price $p_{ij}$ that solves the following problem:
@@
$0$ $=\frac{d\ln\pi_{i}}{d\ln p_{ij}}+\frac{\phi}{1-\phi}\cdot\frac{GFT_{ij}^{i}}{\pi_{i}}\frac{\pi_{j}}{GFT_{ij}^{j}}\cdot\frac{d\ln\pi_{j}}{d\ln p_{ij}},$ (A.7)
@@
$\pi_{i}=p_{ij}q_{ij}+\sum_{k\neq j}p_{ik}q_{ik}-\theta MC_{i}q_{i},$
where $MC_{i}=k_{i}q_{i}^{\frac{1-\theta}{\theta}}.$ The elasticity of the profit $\pi_{i}$ with respect to $p_{ij}$ can be found as:
$\frac{d\ln\pi_{i}}{d\ln p_{ij}}$ $=\frac{q_{ij}}{\pi_{i}}(\varepsilon_{ij}-1)\left(\mu_{ij}^{\text{oligopoly}}MC_{i}-p_{ij}\right),$ (A.8)
@@
$\tilde{\pi}_{i\left(-j\right)}=\sum_{k\neq j}p_{ik}q_{ik}-\theta\widetilde{MC_{i}q_{i}}$.
@@
$\theta\widetilde{MC_{i}q_{i}}$ $=\theta MC_{i}q_{i}\left(\frac{\widetilde{q_{i}}}{q_{i}}\right)^{\frac{1}{\theta}}.$
$=\theta MC_{i}q_{i}\left(1-x_{ij}\right)^{\frac{1}{\theta}}.$
@@
$=q_{ij}\left(p_{ij}-MC_{i}\mu_{ij}^{\text{oligopsony}}\right),$
@@
$\pi_{j}=p_{j}q_{j}-\varrho MC_{j}q_{j}=\frac{\gamma}{\eta-1}MC_{j}q_{j},$
where $MC_{j}$ is defined in equation (A.3), and $q_{j}$ in equation (A.4). The elasticity of the profit $\pi_{j}$ with respect to $p_{ij}$ can be found as:
$\frac{d\ln\pi_{j}}{d\ln p_{ij}}=(\eta-1)s_{ij}.$ (A.10)
@@
$\tilde{\pi}_{j\left(-i\right)}=\frac{\gamma}{\eta-1}\widetilde{MC_{j}q_{j}}$,
where $\varrho\widetilde{MC_{j}q_{j}}$ denotes the total “outside” cost of importer $j$. The assumptions on technology and demand downstream imply that we can write:
$\widetilde{MC_{j}q_{j}}$ $=\left(1-s_{ij}\right)^{\frac{\eta-1}{\rho-1}}MC_{j}q_{j}.$
@@
$\mu_{ij}:=\frac{p_{ij}}{MC_{i}}=(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}+\omega_{ij}\mu_{ij}^{\text{oligopsony}}$ (A.12)
@@ A.2 Proof of Proposition 2
Proposition 2 characterizes how the bilateral markup $\mu_{ij}$ varies with bilateral market shares. Part I states that a positive relationship between $\mu_{ij}$ and the exporter’s supplier share $s_{ij}$ requires $\phi<1$. Part II states that a negative relationship between $\mu_{ij}$ and the importer’s buyer share $x_{ij}$ can arise only if $\phi>0$ and $\theta<1$.
Proof of Part I
The Nash bargaining solution implies that the bilateral markup satisfies
$\mu_{ij}=(1-\omega_{ij})\mu^{\text{oligopoly}}(s_{ij})+\omega_{ij}\mu^{\text{oligopsony}}(x_{ij}),$ (A.15)
with terms defined above. Differentiating with respect to $s_{ij}$ yields
@@
$\frac{\partial\mu_{ij}}{\partial s_{ij}}=(1-\omega_{ij})\frac{\partial\mu^{\text{oligopoly}}_{ij}}{\partial s_{ij}}+\frac{\partial\omega_{ij}}{\partial s_{ij}}\bigl(\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}\bigr).$ (A.16)
Under the maintained assumption $\eta<\rho$, the oligopoly markup is strictly increasing in supplier share, so $\partial_{s}\mu^{\text{oligopoly}}_{ij}>0$, while the oligopsony markdown is strictly lower than the oligopoly markup, $\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}<0$.
Consider first the case $\phi=1$, which implies $\omega_{ij}=1$ for all $s_{ij}$. Hence, $(1-\omega_{ij})=0$ and $\partial\omega_{ij}/\partial s_{ij}=0$. Both terms in (A.16) therefore vanish, so that
$\frac{\partial\mu_{ij}}{\partial s_{ij}}=0\qquad\text{when }\phi=1.$
Thus, a strictly positive relationship between $\mu_{ij}$ and $s_{ij}$ cannot arise under full importer bargaining power.
Now suppose $\phi<1$. In this case $\omega_{ij}(s_{ij})$ is hump-shaped in $s_{ij}$ with $\omega_{ij}(0)=\omega_{ij}(1)=\phi$, implying that there exists $\bar{s}\in(0,1)$ such that $\partial\omega_{ij}/\partial s_{ij}<0$ for all $s_{ij}>\bar{s}$. For such values of $s_{ij}$, the second term in (A.16) is nonnegative, since it is the product of two negative terms. Together with $(1-\omega_{ij})>0$ and $\partial_{s}\mu^{\text{oligopoly}}_{ij}>0$, this implies
$\frac{\partial\mu_{ij}}{\partial s_{ij}}>0\qquad\text{for all }s_{ij}>\bar{s}.$
For $s_{ij}<\bar{s}$, the sign of $\partial\mu_{ij}/\partial s_{ij}$ depends on the relative strength of the two terms in (A.16). In particular, $\partial\mu_{ij}/\partial s_{ij}>0$ if and only if the direct oligopoly effect dominates the bargaining-weight effect, i.e.,
$(1-\omega_{ij})\frac{\partial\mu^{\text{oligopoly}}_{ij}}{\partial s_{ij}}>\left|\frac{\partial\omega_{ij}}{\partial s_{ij}}\bigl(\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}\bigr)\right|>0.$
We conclude that $\partial\mu_{ij}/\partial s_{ij}>0$ can arise only when $\phi<1$. This establishes Part I of Proposition 2. ∎
Proof of Part II.
From (A.15), the bilateral markup $\mu_{ij}$ depends on the importer’s buyer share $x_{ij}$ only through the oligopsony markdown $\mu^{\text{oligopsony}}_{ij}(x_{ij})$. Differentiating with respect to $x_{ij}$:
$\frac{\partial\mu_{ij}}{\partial x_{ij}}=\omega_{ij}\frac{\partial\mu^{\text{oligopsony}}_{ij}}{\partial x_{ij}}.$
If $\phi=0$, then $\omega_{ij}=0$ for all $x_{ij}$, implying $\partial\mu_{ij}/\partial x_{ij}=0$. Hence, a negative relationship between $\mu_{ij}$ and $x_{ij}$ cannot occur if $\phi=0$.
For $\phi>0$, the sign of $\partial_{x}\mu_{ij}$ is equal to the sign of $\partial_{x}\mu^{\text{oligopsony}}_{ij}$. This is given by:
$\frac{\partial\mu^{\text{oligopsony}}_{ij}}{\partial x_{ij}}=\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}-\theta\bigl[1-(1-x_{ij})^{\frac{1}{\theta}}\bigr]}{x_{ij}^{2}}.$
We establish the following lemma.
Lemma 1.
For any $x\in(0,1)$ and $\theta>0$,
$\operatorname{sign}\!\left(x(1-x)^{\frac{1}{\theta}-1}-\theta\bigl[1-(1-x)^{\frac{1}{\theta}}\bigr]\right)=\operatorname{sign}(\theta-1).$
Proof.
Let $t=(1-x)^{1/\theta}\in(0,1)$. Then the expression in the lemma can be written as
$t^{1-\theta}-t-\theta+t\theta.$
Express the above as a function of $\theta$, $f\left(\theta\right)$. One can show that $f\left(\theta\right)$ is strictly convex in $\theta$ ($\partial^{2}f\left(\theta\right)/\partial\theta^{2}>0$), and $f\left(0\right)=f\left(1\right)=0$. This implies
[…]
with equality at $\theta=1$. ∎
By the lemma, $\partial\mu^{\text{oligopsony}}_{ij}/\partial x_{ij}<0$ if and only if $\theta<1$ since $x_{ij}^{2}>0$ for all $x\in(0,1]$. Since $\omega_{ij}>0$ when $\phi>0$, this implies
$\frac{\partial\mu_{ij}}{\partial x_{ij}}<0\quad\Longleftrightarrow\quad\phi>0\text{ and }\theta<1.$
Therefore, observing a negative relationship between the bilateral markup and the importer’s buyer share implies positive importer bargaining power and decreasing returns to scale, establishing Part II of Proposition 2. ∎
@@ A.3 Proof of Proposition 3
$\ln p_{ij}=\ln\mu_{ij}+\ln MC_{i}+\ln T_{c},$
where $MC_{i}$ and $\mu_{ij}$ are as in equations (2.1) and (2.10), respectively.
@@
where $\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}$ and $\Lambda_{ij}\equiv-\frac{d\ln MC_{i}}{d\ln p_{ij}}$ are the partial markup and cost elasticities, respectively.
@@
$\ln MC_{i}$ $=\ln k_{i}+\frac{1-\theta}{\theta}\ln q_{i}.$
@@
$\Lambda_{ij}\equiv-\frac{d\ln MC_{i}}{d\ln p_{ij}}$ $=\frac{1-\theta}{\theta}\frac{d\ln q_{i}}{d\ln q_{ij}}\left(-\frac{d\ln q_{ij}}{d\ln p_{ij}}\right)$
@@
$={}$ $\left[\frac{\left((1-x_{ij})^{\frac{1-\theta}{\theta}}-\frac{1-\theta}{\theta}x_{ij}(1-x_{ij})^{\frac{1}{\theta}-2}\right)\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]-x_{ij}(1-x_{ij})^{\frac{2}{\theta}-2}}{\theta^{2}\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]^{2}}\right](1-x_{ij})$
@@
$={}$ $\left[\left(\frac{1-\frac{1-\theta}{\theta}\cdot\frac{x_{ij}}{1-x_{ij}}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-\frac{(1-x_{ij})^{\frac{1}{\theta}-1}}{\mu_{ij}^{\text{oligopsony}}\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}\right)(1-x_{ij})^{\frac{1}{\theta}}\right$.